Science and medicine. A plane flies 720 mi against a steady 30-mi/h headwind
and then returns to the same point with the wind. If the entire trip takes 10 h, what is the plane’s speed in still air?

Answer by scott8148(5891)

let x=speed ___ time=distance/rate ___ total time is 10

(720/(x-30))+(720/(x+30))=10 ___ (720(x+30))+(720(x-30))=10(x^2-30^2)

x^2-144x-900=0 ___ (x-150)(x+6)=0